40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 15 15:03:41 2022, Koji, Update, PSL, PMC input beam aligned Screenshot_2022-06-15_15-03-16.pngPXL_20220615_212304650.jpg
    Reply  Thu Jun 16 15:29:03 2022, yuta, Update, PSL, PMC input beam aligned again, IMC  Screenshot_2022-06-16_15-27-30.png
       Reply  Wed Nov 16 11:56:21 2022, Radhika, Update, PSL, PMC input beam aligned again, IMC  Screen_Shot_2022-11-16_at_12.04.18_PM.png
          Reply  Mon Nov 21 09:13:31 2022, JC, Update, PSL, PMC input beam aligned again, IMC  
             Reply  Tue Nov 22 08:56:27 2022, JC, Update, PSL, PMC input beam alignment Screenshot_from_2022-11-22_08-58-45.png
                Reply  Mon Nov 28 11:21:25 2022, JC, Update, PSL, PMC input beam alignment Screenshot_from_2022-11-28_11-20-31.png
Message ID: 16922     Entry time: Thu Jun 16 15:29:03 2022     In reply to: 16917     Reply to this: 17271
Author: yuta 
Type: Update 
Category: PSL 
Subject: PMC input beam aligned again, IMC  

[Paco, Tomislav, Yuta]

Somehow, when we were trying to measure WFS open loop transfer functions, PMC unlocked many times for the past two hours and PMC transmission got low.
PMC iput beam was aligned again, and IMC WFS DC offsets and RF offsets were adjusted.
PMC transmission is now C1:PSL-PMC_PMCTRANSPD~0.75, and IMC transmission is C1:IOO-MC_TRANS_SUM~1.4e4.
Actually, IMC transmission once reached 1.5e4 at 06-16-2022 20:01 UTC with PMC transmission of 0.75 (see Attached). There might be a better alignment.

Attachment 1: Screenshot_2022-06-16_15-27-30.png  88 kB  Uploaded Thu Jun 16 16:39:31 2022  | Show | Show all
ELOG V3.1.3-